OpenStudy (anonymous):

Calculate the convergence integration values |cos x|/x^2 dx, from zero to infinity

OpenStudy (amistre64):

do you recall the power series for cos x? that might make life simpler ... and it might not

OpenStudy (amistre64):

\[cos x = 1+\frac{x^2}{2!}+\frac{x^4}{4!}+\frac{x^6}{6!}+...\] \[\frac{cos x}{x^2} = x^{-2}+\frac{x^0}{2!}+\frac{x^2}{4!}+\frac{x^4}{6!}+\frac{x^6}{8!}...\] integrating gives us \[\int \frac{cos x}{x^2}dx = -x^{-1}+\frac{x^1}{2!}+\frac{x^3}{3~4!}+\frac{x^5}{5~6!}+\frac{x^7}{7~8!}...\] at 0 its pointless; so the limit of this as x to inf, just make sure i didnt mess it up along the way ....